A chemist mixes 12 L of a solution that is 45% acid with 8 L of a solution that is 75 acid. What is the percent of acid of the mixture?
Answer by lwsshak3(11628)

A chemist mixes 12 L of a solution that is 45% acid with 8 L of a solution that is 75 acid. What is the percent of acid of the mixture.
..
let x=percent of acid in mixture
liters of mixture=8+12=20
.45*12+.75*8=x*20
5.4+6=20x
x=11.4/20=.57
ans:
The percent of acid of the mixture=57%